# Can AI-Generated Persuasion Be Detected? Persuaficial Benchmark and AI vs. Human Linguistic Differences

### v1

@@ Appendix I Cohen’s d statistic definition and our computation
GPT 4.1 Mini - Paraphrasing
L_CONT_T 0.5054
L_CONT_A 0.4590
SY_INV_PATTERNS -0.4542
LTOKEN_RATIO_LEM 0.3713
L_FUNC_A -0.3619
L_PUNCT_DASH 0.3378
ST_REPET_WORDS -0.3280
L_PUNCT_SEMC -0.2656
L_PUNCT_COM 0.2298
ASM -0.2196
VT_MIGHT 0.1995
L_LINKS -0.1954
L_ADJ_POSITIVE 0.1723
CDS -0.1719
PS_AGREEMENT -0.1667
L_ADV_SUPERLATIVE 0.1664
L_ADV_COMPARATIVE 0.1592
POS_ADV 0.1583
PS_CAUSE -0.1537
PS_TIME -0.1532
@@
L_CONT_T 0.7036
L_CONT_A 0.6222
VT_MIGHT 0.6188
SY_INV_PATTERNS -0.6072
L_PLURAL_NOUNS 0.5444
L_FUNC_A -0.5293
SY_NARRATIVE 0.5242
LTOKEN_RATIO_LEM 0.4993
ST_REPET_WORDS -0.4319
POS_PRO -0.4307
L_YOU_PRON -0.4245
G_ACTIVE -0.4232
G_FUTURE -0.4015
L_SECOND_PERSON_PRON -0.3997
VT_FUTURE_SIMPLE -0.3949
CDS -0.3837
SENT_D_PP 0.3689
L_PUNCT -0.3401
VT_MAY 0.3378
SENT_ST_DIFFERENCE -0.3208
@@
L_CONT_T 0.7685
L_CONT_A 0.7654
L_PUNCT_DASH 0.7243
L_FUNC_A -0.5702
LTOKEN_RATIO_LEM 0.5075
L_ADV_SUPERLATIVE 0.5033
L_ADV_COMPARATIVE 0.4892
SY_INV_PATTERNS -0.4888
POS_ADV 0.4766
ST_REPET_WORDS -0.4037
L_ADJ_POSITIVE 0.3884
POS_ADJ 0.3670
ASM -0.3241
L_ADV_POSITIVE 0.3237
PS_CAUSE -0.3067
L_PUNCT_COM 0.3045
L_PUNCT_SEMC -0.2626
POS_PRO -0.2415
SENT_D_ADVP 0.2385
L_NOUN_PHRASES 0.2333
@@
L_CONT_T 1.6204
L_CONT_A 1.4300
LTOKEN_RATIO_LEM 1.3172
SENT_D_NP 1.0532
ST_REPET_WORDS -0.9950
L_PUNCT_DASH 0.9885
L_FUNC_A -0.8585
SY_IMPERATIVE 0.8376
POS_NOUN 0.8076
POS_ADJ 0.7841
L_ADJ_POSITIVE 0.7816
SY_INV_PATTERNS -0.7496
L_LINKS -0.7474
VF_INFINITIVE 0.7252
G_PAST -0.6604
G_ACTIVE -0.6565
L_FIRST_PERSON_SING_PRON -0.6425
L_I_PRON -0.6425
SENT_D_VP 0.6275
VT_PAST_SIMPLE -0.5774
@@
SENT_ST_WPERSENT 0.7055
SENT_ST_DIFFERENCE -0.6979
L_PUNCT_COM 0.6659
SY_INV_PATTERNS -0.6146
L_CONT_T 0.5438
L_CONT_A 0.4861
G_ACTIVE -0.4821
L_PUNCT_DOT -0.4509
ASM -0.4192
POS_PRO -0.4055
FOS_FRONTING 0.3884
L_ADJ_POSITIVE 0.3844
L_PUNCT_SEMC -0.3839
L_YOU_PRON -0.3761
L_FUNC_A -0.3469
SY_SUBORD_SENT 0.3354
L_PUNCT -0.3339
POS_PREP 0.3256
VT_PRESENT_SIMPLE -0.3229
L_SECOND_PERSON_PRON -0.3228
@@
SY_INV_PATTERNS -0.7470
G_ACTIVE -0.7152
SENT_ST_WPERSENT 0.6961
L_CONT_T 0.6226
SENT_D_PP 0.6153
L_ADJ_POSITIVE 0.6145
POS_PREP 0.6086
L_YOU_PRON -0.5966
POS_NOUN 0.5896
L_PUNCT_COM 0.5848
L_SECOND_PERSON_PRON -0.5764
POS_ADJ 0.5595
SY_NARRATIVE 0.5419
L_CONT_A 0.5413
POS_PRO -0.5400
SENT_ST_DIFFERENCE -0.5382
SY_SUBORD_SENT 0.5373
L_PUNCT -0.5195
ASM -0.4923
L_PLURAL_NOUNS 0.4688
@@
SENT_ST_WPERSENT 0.8019
L_ADJ_POSITIVE 0.7196
L_PUNCT_COM 0.7055
L_CONT_T 0.6589
POS_ADJ 0.6478
L_CONT_A 0.6349
G_ACTIVE -0.6171
SENT_ST_DIFFERENCE -0.6110
L_PUNCT_DOT -0.5876
SY_INV_PATTERNS -0.5135
ASM -0.4697
L_FUNC_A -0.4692
L_FUNC_T -0.4456
FOS_FRONTING 0.4390
POS_PRO -0.3973
SENT_D_VP 0.3895
L_PUNCT_SEMC -0.3806
CDS -0.3720
L_YOU_PRON -0.3670
L_NOUN_PHRASES 0.3630
@@
VF_INFINITIVE 1.3397
SY_IMPERATIVE 1.2406
G_ACTIVE -0.9149
SENT_D_VP 0.8331
L_PUNCT_DOT -0.8088
G_PAST -0.7696
L_OUR_PRON 0.7642
L_LINKS -0.7474
VT_PAST_SIMPLE -0.6985
SY_EXCLAMATION 0.6879
SY_INV_PATTERNS -0.6857
L_CONT_T 0.6780
SY_COORD_SENT 0.6774
L_PUNCT -0.6493
SENT_D_NP 0.6396
L_FIRST_PERSON_SING_PRON -0.6230
L_I_PRON -0.6230
L_WE_PRON 0.6190
L_IT_PRON 0.5548
VT_MUST 0.5078
@@
L_CONT_T 0.7526
L_CONT_A 0.6659
L_FUNC_A -0.6181
SY_INV_PATTERNS -0.6133
L_ADV_SUPERLATIVE 0.4274
ST_REPET_WORDS -0.4178
L_ADV_COMPARATIVE 0.4141
LTOKEN_RATIO_LEM 0.4081
PS_CONDITION -0.3995
ASM -0.3761
CDS -0.3606
POS_ADV 0.3605
L_ADJ_POSITIVE 0.3407
PS_AGREEMENT -0.3212
L_PUNCT_COM 0.3172
PS_CAUSE -0.2988
POS_ADJ 0.2872
POS_PRO -0.2805
G_ACTIVE -0.2710
SENT_ST_WPERSENT 0.2582
@@
L_CONT_T 1.0245
L_CONT_A 1.0176
L_FUNC_A -0.8850
POS_PRO -0.7267
SY_INV_PATTERNS -0.7119
L_PLURAL_NOUNS 0.6762
POS_NOUN 0.6717
SENT_D_PP 0.6401
L_ADJ_POSITIVE 0.6191
G_ACTIVE -0.5837
L_SECOND_PERSON_PRON -0.5783
L_YOU_PRON -0.5736
SY_NARRATIVE 0.5666
POS_ADJ 0.5620
CDS -0.5438
L_FUNC_T -0.4852
VF_INFINITIVE -0.4846
ASM -0.4624
L_THEY_PRON -0.4486
POS_CONJ -0.4429
@@
L_CONT_T 1.0424
L_CONT_A 1.0122
L_FUNC_A -0.9718
L_ADJ_POSITIVE 0.9024
POS_ADJ 0.8267
PS_CONDITION -0.6942
PS_CAUSE -0.6071
L_FUNC_T -0.6021
SY_INV_PATTERNS -0.5923
G_ACTIVE -0.5882
L_NOUN_PHRASES 0.5486
POS_PRO -0.5389
L_ADV_SUPERLATIVE 0.5243
L_ADV_COMPARATIVE 0.5002
ASM -0.4781
CDS -0.4777
L_SINGULAR_NOUNS 0.4707
POS_NOUN 0.4570
L_YOU_PRON -0.4543
SENT_D_NP 0.4475
@@
SY_IMPERATIVE 1.7354
L_CONT_T 1.4021
L_CONT_A 1.2523
SENT_D_NP 1.1601
PS_CAUSE -1.1529
VF_INFINITIVE 1.0451
L_FUNC_A -0.9623
PS_CONDITION -0.8951
POS_NOUN 0.8730
ST_REPET_WORDS -0.8055
LTOKEN_RATIO_LEM 0.7851
L_LINKS -0.7474
SY_INV_PATTERNS -0.7306
L_SINGULAR_NOUNS 0.6881
SY_EXCLAMATION 0.6825
SENT_D_VP 0.6631
POS_PREP -0.6571
L_I_PRON -0.6461
L_FIRST_PERSON_SING_PRON -0.6461
L_NOUN_PHRASES 0.6320
@@
L_CONT_A 0.7275
L_CONT_T 0.7074
SY_INV_PATTERNS -0.5802
L_FUNC_A -0.5371
L_PUNCT_COM 0.4880
LTOKEN_RATIO_LEM 0.4192
L_ADJ_POSITIVE 0.4104
ST_REPET_WORDS -0.3981
POS_ADJ 0.3677
L_NOUN_PHRASES 0.3255
ASM -0.2996
PS_CONDITION -0.2850
PS_CAUSE -0.2759
L_POSSESSIVES 0.2719
POS_PREP -0.2305
CDS -0.2242
L_PUNCT 0.2218
PS_AGREEMENT -0.2169
L_PLURAL_NOUNS 0.2053
L_THEIR_PRON 0.2050
@@
L_CONT_A 0.9216
L_CONT_T 0.9163
L_FUNC_A -0.7257
SY_INV_PATTERNS -0.7182
POS_NOUN 0.5946
G_ACTIVE -0.5892
POS_PRO -0.5535
VT_MIGHT 0.5222
CDS -0.5034
SENT_D_PP 0.4976
L_YOU_PRON -0.4908
LTOKEN_RATIO_LEM 0.4813
L_PLURAL_NOUNS 0.4761
L_ADJ_POSITIVE 0.4695
L_SECOND_PERSON_PRON -0.4555
ASM -0.4517
L_PUNCT_COM 0.4347
ST_REPET_WORDS -0.4182
POS_ADJ 0.4165
SY_NARRATIVE 0.4079
@@
L_CONT_A 1.0170
L_CONT_T 0.9948
L_FUNC_A -0.8186
L_ADJ_POSITIVE 0.7827
POS_ADJ 0.7239
L_PUNCT_COM 0.5901
L_NOUN_PHRASES 0.5756
SY_INV_PATTERNS -0.5355
LTOKEN_RATIO_LEM 0.4892
L_ADV_SUPERLATIVE 0.4824
PS_CONDITION -0.4649
L_ADV_COMPARATIVE 0.4629
ASM -0.4428
ST_REPET_WORDS -0.4386
PS_CAUSE -0.4330
POS_ADV 0.4270
G_ACTIVE -0.4252
POS_PRO -0.3773
L_FUNC_T -0.3605
PS_AGREEMENT -0.3525
@@
L_CONT_T 1.5669
L_CONT_A 1.3747
SY_IMPERATIVE 1.3186
LTOKEN_RATIO_LEM 1.1953
SY_EXCLAMATION 1.1092
VF_INFINITIVE 1.0846
ST_REPET_WORDS -1.0032
SENT_D_VP 0.9291
L_FUNC_A -0.8914
POS_NOUN 0.8779
L_NOUN_PHRASES 0.8190
PS_CAUSE -0.7878
PS_CONDITION -0.7832
G_ACTIVE -0.7769
L_LINKS -0.7474
L_SINGULAR_NOUNS 0.7331
SY_INV_PATTERNS -0.6693
G_PAST -0.6464
POS_PREP -0.6279
L_I_PRON -0.6236

### v2

@@ Appendix I Statistical Analysis with Wilcoxon Signed-rank Test
GPT-4.1-mini - Paraphrased
L_TYPE_TOKEN_RATIO_LEMMAS 0.9218
L_CONT_T 0.8239
ST_REPETITIONS_WORDS -0.7487
L_CONT_A 0.6522
L_FUNC_A -0.5824
SY_INVERSE_PATTERNS -0.4217
L_PUNCT_COM 0.3217
L_PUNCT_DASH 0.3116
SENT_D_NP 0.3055
L_LINKS -0.2848
L_PLURAL_NOUNS 0.2717
L_ADJ_POSITIVE 0.2461
SENT_ST_DIFFERENCE -0.2374
POS_PRO -0.2357
L_YOU_PRON -0.2195
L_SECOND_PERSON_PRON -0.2118
POS_ADJ 0.2080
L_PUNCT_SEMC -0.1991
ASM -0.1920
L_ADV_SUPERLATIVE 0.1919
@@
L_TYPE_TOKEN_RATIO_LEMMAS 0.9407
L_CONT_T 0.8198
ST_REPETITIONS_WORDS -0.7602
L_PLURAL_NOUNS 0.6603
L_CONT_A 0.6331
L_FUNC_A -0.6062
POS_PRO -0.5107
SY_INVERSE_PATTERNS -0.4739
SY_NARRATIVE 0.4704
L_SECOND_PERSON_PRON -0.4526
VT_MIGHT 0.4523
SENT_D_PP 0.4265
G_ACTIVE -0.4122
SENT_ST_DIFFERENCE -0.4109
L_YOU_PRON -0.3956
L_I_PRON -0.3724
L_FIRST_PERSON_SING_PRON -0.3724
CDS -0.3685
POS_NOUN 0.3636
G_FUTURE -0.3284
@@
L_CONT_T 1.1821
L_TYPE_TOKEN_RATIO_LEMMAS 1.0916
L_CONT_A 1.0240
L_FUNC_A -0.8593
ST_REPETITIONS_WORDS -0.8215
L_PUNCT_DASH 0.5816
L_ADJ_POSITIVE 0.5443
POS_ADJ 0.5200
L_ADV_SUPERLATIVE 0.4963
POS_ADV 0.4924
L_ADV_COMPARATIVE 0.4869
SY_INVERSE_PATTERNS -0.4506
POS_PRO -0.4148
L_ADV_POSITIVE 0.3644
L_PUNCT_COM 0.3578
SENT_D_NP 0.3522
L_NOUN_PHRASES 0.3414
PS_CAUSE -0.3385
L_LINKS -0.3305
SENT_ST_WRDSPERSENT 0.3042
@@
L_CONT_T 1.4059
L_CONT_A 1.2264
L_TYPE_TOKEN_RATIO_LEMMAS 1.0753
ST_REPETITIONS_WORDS -0.8541
SENT_D_NP 0.7600
L_FUNC_A -0.7259
POS_NOUN 0.7240
L_PUNCT_DASH 0.7067
POS_ADJ 0.6430
L_ADJ_POSITIVE 0.6377
SY_IMPERATIVE 0.5960
VF_INFINITIVE 0.5668
G_PAST -0.5589
SY_INVERSE_PATTERNS -0.5410
L_LINKS -0.5285
G_ACTIVE -0.5023
SENT_D_VP 0.4973
VT_PAST_SIMPLE -0.4660
L_SINGULAR_NOUNS 0.4653
L_I_PRON -0.4632
@@
SENT_ST_WRDSPERSENT 0.7721
SENT_ST_DIFFERENCE -0.7695
L_CONT_T 0.7592
L_PUNCT_COM 0.6832
L_TYPE_TOKEN_RATIO_LEMMAS 0.6077
L_CONT_A 0.5676
L_PUNCT_DOT -0.5399
POS_PRO -0.5077
SY_INVERSE_PATTERNS -0.4720
G_ACTIVE -0.4647
L_FUNC_A -0.4470
L_SECOND_PERSON_PRON -0.4128
L_YOU_PRON -0.4097
ST_REPETITIONS_WORDS -0.3859
L_ADJ_POSITIVE 0.3841
VT_PRESENT_SIMPLE -0.3681
L_PROPER_NAME -0.3615
L_LINKS -0.3494
POS_NOUN 0.3463
ASM -0.3324
@@
SENT_ST_WRDSPERSENT 0.7274
L_CONT_T 0.7048
SENT_D_PP 0.6735
SENT_ST_DIFFERENCE -0.6391
L_PROPER_NAME -0.6114
POS_NOUN 0.6002
G_ACTIVE -0.5923
SY_INVERSE_PATTERNS -0.5443
L_ADJ_POSITIVE 0.5337
L_PUNCT_COM 0.5158
POS_PRO -0.5124
L_CONT_A 0.5080
L_SECOND_PERSON_PRON -0.5064
L_PLURAL_NOUNS 0.5021
POS_PREP 0.4996
POS_ADJ 0.4885
L_TYPE_TOKEN_RATIO_LEMMAS 0.4744
L_YOU_PRON -0.4730
SY_NARRATIVE 0.4673
SY_SUBORD_SENT 0.4254
@@
L_CONT_T 0.9518
SENT_ST_WRDSPERSENT 0.8794
L_CONT_A 0.7635
L_ADJ_POSITIVE 0.7466
SENT_ST_DIFFERENCE -0.7274
L_PUNCT_COM 0.7113
POS_ADJ 0.6743
L_FUNC_T -0.6440
G_ACTIVE -0.6295
L_FUNC_A -0.6168
L_PUNCT_DOT -0.5948
L_PROPER_NAME -0.5737
L_TYPE_TOKEN_RATIO_LEMMAS 0.4713
POS_PRO -0.4696
G_PAST -0.4403
SY_INVERSE_PATTERNS -0.4362
L_SECOND_PERSON_PRON -0.4237
L_NOUN_PHRASES 0.4208
L_YOU_PRON -0.4169
SENT_D_VP 0.4091
@@
VF_INFINITIVE 1.0143
SY_IMPERATIVE 0.8892
G_ACTIVE -0.6972
SENT_D_VP 0.6751
L_OUR_PRON 0.6048
G_PAST -0.6043
L_CONT_T 0.5952
L_PUNCT_DOT -0.5939
VT_PAST_SIMPLE -0.5411
L_LINKS -0.5285
SY_EXCLAMATION 0.5258
SY_COORD_SENT 0.5123
SY_INVERSE_PATTERNS -0.5000
L_PUNCT -0.4763
SENT_D_NP 0.4762
L_WE_PRON 0.4659
L_I_PRON -0.4463
L_FIRST_PERSON_SING_PRON -0.4463
POS_NOUN 0.4267
L_PERSONAL_NAME -0.4258
@@
L_CONT_T 0.8744
L_TYPE_TOKEN_RATIO_LEMMAS 0.7274
L_FUNC_A -0.6963
L_CONT_A 0.6769
ST_REPETITIONS_WORDS -0.6665
SY_INVERSE_PATTERNS -0.4798
L_ADV_SUPERLATIVE 0.3887
L_ADV_COMPARATIVE 0.3845
SENT_ST_WRDSPERSENT 0.3600
PS_CONDITION -0.3599
POS_PRO -0.3556
POS_ADV 0.3344
L_PUNCT_COM 0.3338
CDS -0.3291
L_ADJ_POSITIVE 0.3191
L_SECOND_PERSON_PRON -0.3096
SENT_ST_DIFFERENCE -0.3074
L_FUNC_T -0.3063
L_YOU_PRON -0.2964
ASM -0.2902
@@
L_CONT_T 1.0167
L_CONT_A 0.8716
L_TYPE_TOKEN_RATIO_LEMMAS 0.8497
L_FUNC_A -0.7555
ST_REPETITIONS_WORDS -0.6725
POS_NOUN 0.6419
POS_PRO -0.5732
SY_INVERSE_PATTERNS -0.5399
L_PLURAL_NOUNS 0.5274
G_ACTIVE -0.5179
SENT_D_PP 0.5060
L_SECOND_PERSON_PRON -0.4606
SENT_D_NP 0.4539
CDS -0.4498
L_YOU_PRON -0.4263
L_ADJ_POSITIVE 0.4165
L_PUNCT_COM 0.3967
L_NOUN_PHRASES 0.3906
SY_NARRATIVE 0.3875
VT_MIGHT 0.3785
@@
L_CONT_T 1.1099
L_FUNC_A -0.9504
L_CONT_A 0.9499
L_ADJ_POSITIVE 0.8330
POS_ADJ 0.7743
L_FUNC_T -0.7336
L_TYPE_TOKEN_RATIO_LEMMAS 0.5909
ST_REPETITIONS_WORDS -0.5650
PS_CONDITION -0.5605
SENT_D_NP 0.5390
G_ACTIVE -0.5361
POS_PRO -0.5318
L_NOUN_PHRASES 0.5293
PS_CAUSE -0.5251
L_PROPER_NAME -0.4946
CDS -0.4634
L_SINGULAR_NOUNS 0.4587
POS_NOUN 0.4537
SY_INVERSE_PATTERNS -0.4488
L_ADV_SUPERLATIVE 0.4420
@@
SY_IMPERATIVE 1.2484
L_CONT_T 1.2073
L_CONT_A 1.0548
L_FUNC_A -0.9226
PS_CAUSE -0.8919
VF_INFINITIVE 0.8048
SENT_D_NP 0.8033
PS_CONDITION -0.6801
POS_NOUN 0.6799
ST_REPETITIONS_WORDS -0.6654
L_TYPE_TOKEN_RATIO_LEMMAS 0.6102
L_NOUN_PHRASES 0.5410
L_LINKS -0.5285
L_SINGULAR_NOUNS 0.5261
SY_INVERSE_PATTERNS -0.5239
SY_EXCLAMATION 0.5174
SENT_D_VP 0.5048
G_PAST -0.5028
L_HASHTAG 0.4954
POS_PREP -0.4900
@@
L_CONT_T 0.8961
L_TYPE_TOKEN_RATIO_LEMMAS 0.8198
L_CONT_A 0.7892
ST_REPETITIONS_WORDS -0.7241
L_FUNC_A -0.6619
L_PUNCT_COM 0.5714
SY_INVERSE_PATTERNS -0.4691
L_ADJ_POSITIVE 0.4200
POS_ADJ 0.3799
L_NOUN_PHRASES 0.3709
SENT_D_NP 0.3249
POS_PRO -0.2846
PS_CONDITION -0.2802
L_POSSESSIVES 0.2743
L_PLURAL_NOUNS 0.2732
L_PUNCT 0.2724
PS_CAUSE -0.2699
L_PUNCT_DOT 0.2668
VF_INFINITIVE -0.2396
ASM -0.2375
@@
L_CONT_T 1.0167
L_CONT_A 0.8716
L_TYPE_TOKEN_RATIO_LEMMAS 0.8497
L_FUNC_A -0.7555
ST_REPETITIONS_WORDS -0.6725
POS_NOUN 0.6419
POS_PRO -0.5732
SY_INVERSE_PATTERNS -0.5399
L_PLURAL_NOUNS 0.5274
G_ACTIVE -0.5179
SENT_D_PP 0.5060
L_SECOND_PERSON_PRON -0.4606
SENT_D_NP 0.4539
CDS -0.4498
L_YOU_PRON -0.4263
L_ADJ_POSITIVE 0.4165
L_PUNCT_COM 0.3967
L_NOUN_PHRASES 0.3906
SY_NARRATIVE 0.3875
VT_MIGHT 0.3785
@@
L_CONT_T 1.2269
L_CONT_A 1.1194
L_FUNC_A -0.9824
L_TYPE_TOKEN_RATIO_LEMMAS 0.8932
L_ADJ_POSITIVE 0.8740
POS_ADJ 0.8168
ST_REPETITIONS_WORDS -0.7448
L_NOUN_PHRASES 0.6731
L_PUNCT_COM 0.6116
L_FUNC_T -0.5222
POS_PRO -0.4680
G_ACTIVE -0.4623
PS_CAUSE -0.4416
L_ADV_SUPERLATIVE 0.4368
PS_CONDITION -0.4344
SY_INVERSE_PATTERNS -0.4317
SENT_D_NP 0.4279
L_ADV_COMPARATIVE 0.4232
POS_NOUN 0.4109
G_PAST -0.4060
@@
L_CONT_T 1.3409
L_CONT_A 1.1627
L_TYPE_TOKEN_RATIO_LEMMAS 0.9696
SY_IMPERATIVE 0.9341
ST_REPETITIONS_WORDS -0.8302
VF_INFINITIVE 0.8296
SY_EXCLAMATION 0.8239
L_FUNC_A -0.8082
POS_NOUN 0.6966
L_NOUN_PHRASES 0.6962
SENT_D_VP 0.6846
PS_CAUSE -0.6017
PS_CONDITION -0.5881
L_SINGULAR_NOUNS 0.5842
G_ACTIVE -0.5623
L_HASHTAG 0.5397
L_LINKS -0.5285
G_PAST -0.5183
SY_INVERSE_PATTERNS -0.4737
POS_PREP -0.4627